The generator at a power plant produces AC at 25,000 V. A transformer steps this up to 355,000 V for transmission over power lines. If there are 2,150 turns of wire in the input coil of the transformer, how many turns must there be in the output coil?
1
Expert's answer
2019-05-06T10:36:29-0400

According to the ideal transformer identity,


"\frac{V_P}{V_S} = \frac{N_P}{N_S},"

where indices "P" and "S" correspond to the "primary" (input) and "secondary" (output) parameters.

Hence,


"N_S = N_P \frac{V_S}{V_P}"

Substituting the numerical values, we obtain:


"N_S = 2150 \cdot \frac{355 \cdot 10^{3}}{25 \cdot 10^{3}} = 30530"

Answer: 30530 turns.